# CLARA: A Modular Framework for Unsupervised Transit Detection Using TESS Light Curves

### v1

@@ Acknowledgements
Questions regarding the dataset or reproducibility materials should be directed to the corresponding author via email or github issues.
Appendix A Notes
A.1 URF-3 Cross-Sector Validation
URF-3’s performance exhibits significant sector-dependent variation due to its training methodology. Since URF-3 uses feature vectors extracted from confirmed TOIs in Sector 1 as its synthetic anomaly class, it demonstrates clear overfitting when evaluated on the same sector. This is evident in the score distribution plots (Figures 2 and 3), where URF-3’s anomaly scores for TOIs in Sector 1 are systematically higher (peaking around 0.45-0.47) compared to Sector 2 (peaking around 0.10-0.12). The reduced score separation in Sector 2 reflects URF-3’s diminished ability to generalize beyond the specific feature characteristics of its training TOIs. This sector dependency underscores the importance of independent validation sets and highlights a fundamental limitation of using confirmed target features directly as synthetic training data, reinforcing URF-4’s advantage in using parametric synthetic generation for improved generalization.
A.2 Hyper-parameter Search Space
@@
A.3 Note on Duration Units
@@ A.4 Note on TOI Importance AUC
A.5 Note on Model Selection Strategy: Minimal Supervision Framework
@@
Appendix B Astrometric and Stellar Features Description
@@
Appendix C Tables and Data
Table 8: Comparison of detection efficiency and enrichment of various transit detection pipelines. Detection rate is defined as the percentage of transit-like signals (confirmed or candidate) identified among the total light curves analyzed.
Method Light Curves per Sector Total Light Curves Project Candidates Detection Rate
CLARA Pipeline (till DTW-based DBSCAN clustering) $\sim 90$ 453 19 4.20%
DPMM Cluster 2 $\sim 23$ 114 16 14.04%
TESS SPOC (All Sectors, 2025) $\sim 20,000$ 1,680,000 (84 sectors) 7658 0.4569%
TESS-SPOC FFI Targets (Doyle et al., 2024) 2,890,583 7658 0.2650%
Appendix D Candidates from SPOC sector 1

### v3

@@ 6 Data Availability
This research was conducted entirely independently and without institutional support or funding. All computations were performed on a personal desktop machine (Intel i3 processor, 32 GB RAM), which, despite its modest capacity, was sufficient for this work. The author thanks the developers and maintainers of the open-source software and data archives acknowledged in the Data Availability section and Computational Resources (A)section, which made this work possible. {contribution}Mainak Dasgupta conceived the research problem, designed and implemented the CLARA methodology, performed the simulations and analysis, and wrote the manuscript. The codebase and documentation were also developed and maintained by the author.
Appendix A Computational Resources and Runtime
Appendix B Notes
B.1 URF-3 Cross-Sector Validation
URF-3’s performance exhibits significant sector-dependent variation due to its training methodology. Since URF-3 uses feature vectors extracted from confirmed TOIs in Sector 1 as its synthetic anomaly class, it demonstrates clear overfitting when evaluated on the same sector. URF-3’s anomaly scores for TOIs in Sector 1 are systematically higher (peaking around 0.45-0.47 compared to a Non-TOI distribution peaking at 0.10-0.12) compared to Sector 2 (peaking around 0.10-0.12). The reduced score separation in Sector 2 reflects URF-3’s diminished ability to generalize beyond the specific feature characteristics of its training TOIs. This sector dependency underscores the importance of independent validation sets and highlights a fundamental limitation of using confirmed target features directly as synthetic training data, reinforcing URF-4’s advantage in using parametric synthetic generation for improved generalization.
B.2 Hyper-parameter Search Space
@@
B.3 Note on Duration Units
@@ B.4 Note on TOI Importance AUC
B.5 Note on Model Selection Strategy: Minimal Supervision Framework
@@
Appendix C Astrometric and Stellar Features Description
@@
Appendix D Tables and Figures
Variant TP FP TP+FP TN Precision Recall
URF-1 22 2159 2181 0 1.01 12.57
URF-2 45 3990 4035 0 1.12 25.71
URF-3 14 300 314 15499 4.46 7.18
URF-4 64 4982 5046 0 1.27 36.57
Table 11: Initial injection–recovery test results at threshold = 0.35. Each configuration consisted of 20 injected and 1980 real TESS light curves.
Config WRSS Retrieved Curves to Vet Recovery Ratio
C1 20 119 0.168
C2 20 125 0.160
C3 20 121 0.165
C4 20 122 0.164
C5 20 120 0.167
C6 20 130 0.154
C7 19 124 0.153
C8 19 136 0.140
C9 19 120 0.158
C10 19 142 0.134
Table 12: Initial injection–recovery test results at threshold = 0.60. Each configuration consisted of 20 injected and 1980 real TESS light curves.
Config WRSS Retrieved Curves to Vet Recovery Ratio
C1 17 22 0.773
C2 16 24 0.667
C3 19 24 0.792
C4 18 25 0.720
C5 18 23 0.783
C6 14 23 0.609
C7 16 23 0.696
C8 17 27 0.630
C9 16 21 0.762
C10 16 31 0.516
Table 13: Initial injection–recovery test results at threshold = 0.90. Each configuration consisted of 20 injected and 1980 real TESS light curves.
Config WRSS Retrieved Curves to Vet Recovery Ratio
C1 8 8 1.00
C2 4 4 1.00
C3 8 8 1.00
C4 4 4 1.00
C5 10 10 1.00
C6 2 2 1.00
C7 6 6 1.00
C8 4 4 1.00
C9 3 3 1.00
C10 4 4 1.00
Figure 18: ncurves, duration_hr (days B.3), noise (ppm) and cadence (minutes) clockwise from top left against TOI Recall (blue) and Anomaly Rate (red)
Figure 19: Synthetic features (ncurves, duration_hr(in days B.3), noise (ppm), cadence (minutes) - clockwise from top left) v.s TOI Importance AUC
@@ Appendix D Tables and Figures
Table 16: TOIs in DPMM Cluster 2
# TIC ID TOI
1 92352620 107.01
2 144065872 105.01
3 149603524 102.01
4 260609205 219.01
5 38846515 106.01
6 149603524 102.01
7 184240683 250.01
8 230982885 195.01
9 388104525 112.01
10 38846515 106.01
11 260609205 219.01
12 349518800 407.01
13 388104525 112.01
14 122612091 264.01
15 38686737 432.01
16 38846515 106.01
Table 17: TOIs in GMM Cluster 0
# TIC ID
1 92352620
2 144065872
3 149603524
4 260609205
5 38846515
6 149603524
7 184240683
8 230982885
9 388104525
10 38846515
11 100100827
12 260609205
13 349518800
14 388104525
15 122612091
16 38686737
17 38846515
Table 18: Configuration ranges for the pool of 200 synthetic injected transit curves. Each configuration contributes 20 curves, with depth ranges expressed in parts per million (ppm) and white noise levels specified in ppm.
Config ID Depth Range (ppm) Noise (ppm) # Curves
C1 200–500 50 20
C2 200–500 200 20
C3 500–1000 50 20
C4 500–1000 300 20
C5 1000–2000 100 20
C6 1000–2000 400 20
C7 2000–5000 150 20
C8 2000–5000 500 20
C9 5000–10000 300 20
C10 10000–20000 600 20
Figure 24: Monte Carlo Injection-Recovery performance metrics. Each panel shows the mean (solid line) and ±1 standard deviation (shaded area) for (a) Precision, (b) Recall, (c) Number of injected signals retrieved, and (d) Total number of signals retrieved. The results are based on Monte Carlo injection-recovery tests, illustrating how threshold choice affects detection performance.